# Few-shot Tuning of Foundation Models for Class-incremental Learning

### v1

@@ 4 Experiments
4.1 Baselines
In the absence of available FSCIL with foundation model, we first establish a few baselines to better evaluate our proposed framework. To this end, we adopt three prominent methods on which to build these baselines. First, we use prototype learning, which has been shown to perform well for learning incremental classes in traditional FSCIL [42]. To apply this approach to a foundation model $f_{\theta}(x)$, we keep the encoder frozen during the incremental learning sessions, while the classifier ($W$) is learned with the prototypes [42] of the new classes: $W=\{w_{0}^{0},w_{1}^{0},\cdot\cdot,w_{|c^{0}|}^{0}\}\cup\{w_{0}^{T},w_{1}^{T},\cdot\cdot,w_{|c^{T}|}^{T}\}$. Here, the prototype $w_{c}^{t}$ of class $c$ from session $t$ is defined as the average of the embeddings of all samples of the class: ${w}_{c}^{t}=\frac{1}{|P_{i}|}\sum_{j\in P_{i}}{f_{\theta}({{x}}_{j})}$, where $P_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\{j:y_{j}=i\}$. Next, we consider linear tuning as a baseline and train the model in an incremental-frozen approach, where the encoder is trained only on the first session, followed by prototype learning for the remaining sessions [50]. Finally, we use parameter-efficient fine-tuning as the third baseline, where we add LoRA to different layers of the pre-trained encoder and train newly added LoRA while keeping the encoder frozen.
4.2 Experiment setup
@@
Dataset. We comprehensively evaluate the performance of our proposed method on a diverse set of 16 datasets, including generic object detection (Caltech101 [12], CIFAR-100 [26], CUB-200 [45], MiniImageNet [40], VOC 2007 [11]), fine-grained recognition (OxfordPets [34], StanfordCars [25], Flower102 [33], Food101 [1], FGVCAircraft [29]), scene recognition (SUN397 [46], Country211 [36]), satellite-image (EuroSAT [14], Resisc-45 [5]), texture recognition (DTD [8]), and traffic sign recognition (GTSRB [17]).
Method Acc
Prototype learning 58.77$\pm$0.0
Linear tuning 58.89$\pm$0.23
LoRA 58.96$\pm$0.11
CoACT 61.43$\pm$0.12
Table 1: Overall performance over 16 datasets.
@@
Implementation details. By default, we divide the classes into 10 (or 9; see Appendix for details) sessions with an equal number of classes and perform 10-shot continual training following FSCIL literature [50; 51]. The performance after each session is calculated as the average accuracy of all classes seen so far. Unless specified otherwise, accuracy refers to the accuracy after the last session, which is the average accuracy over all classes. We also evaluate CoACT on the traditional FSCIL setup, where the initial training is performed with a large labelled set from the same distribution as the incremental sessions, followed by tuning on the few-shot incremental sessions. We use a ViT-B/16 as the backbone for most of the experiments, while we present the detailed per-dataset results with ViT-B/32 and ViT-L/16 as well. The encoders are pre-trained on ImgeNet-21K [40]. The framework is implemented on PyTorch, and experiments are conducted on an Nvidia V100 GPU, where the training takes about 6 hours. More details on the hyper-parameters are provided in Appendix A.2.
4.3 Results
Figure 5: (left) Forgetting of learned classes. (right) Accuracy breakdown over the first session, remaining sessions, and all sessions.
Table 1 presents the average accuracy and standard deviation (over 3 runs) of our method over 16 datasets, in comparison to the aforementioned baselines. As we find from the table, CoACT achieves an average accuracy of 61.43% over the 16 datasets, outperforming others by an average of 2.47%. Among the baselines, LoRA shows the second-best accuracy with 58.96%. The other two baselines show relatively similar average accuracies to LoRA. Note that among the baselines, prototype learning with a frozen encoder is a deterministic process, thus a standard deviation of 0. Note that as we will later demonstrate in Table 3, the SOTA in the standard FSCIL setup achieve $\sim$1gains versus prior work.
Table 2: Performance of CoACT on 16 datasets and its comparison to the baseline methods.
@@ 4.3 Results
ViT-L/16 87.25 79.27 7.45 78.29 56.19 66.36 21.24 74.99 17.22 97.18 98.18 87.95 62.25 25.43 67.18 62.23 61.79
[…]
Figure 6: Performance vs. different shots.
Figure 7: Comparison to SOTA on traditional FSCIL.
Figure 8: Forgetting over incr. sessions.
@@
Next, we report the forgetting of learned classes for each method, measured as the drop in accuracy w.r.t. the first session, in Figure 5 (left). As we observe in this figure, CoACT has the least amount of forgetting compared to the baselines with approximately 1.5% less forgetting than both LoRA and linear tuning, and 3.2% than prototype learning. We also present a breakdown of accuracies into the first incremental session, the remaining incremental sessions, and all sessions in Figure 5 (right). Here, for all methods, the first session shows particularly higher accuracy than other sessions since there is no interference (or forgetting) of other classes. While LoRA and linear tuning have higher overall accuracy than prototype tuning, the improvement mainly comes from the higher accuracy in the first session only. All baselines, however, perform relatively similarly in the remaining sessions. In contrast, CoACT shows higher improvement in both the first and remaining sessions.

### v2

@@ 4 Experiments
Table 1: Comparison to prior works across the base and incremental sessions on CUB-200 with 100 base classes and 10-way 5-shot evaluation setup.
Method PT. Backbone Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8 9 10
iCaRL (Rebuffi et al., 2017) 68.68 52.65 48.61 44.16 36.62 29.52 27.83 26.26 24.01 23.89 21.16
EEIL (Castro et al., 2018) 68.68 53.63 47.91 44.20 36.30 27.46 25.93 24.70 23.95 24.13 22.11
TOPIC (Tao et al., 2020) 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28
Rebalancing (Hou et al., 2019) 68.68 57.12 44.21 28.78 26.71 25.66 24.62 21.52 20.12 20.06 19.87
SPPR (Zhu et al., 2021) 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33
MetaFSCIL(Chi et al., 2022) 75.90 72.41 68.78 64.78 62.96 59.99 58.30 56.85 54.78 53.82 52.64
F2M (Shi et al., 2021) 81.07 78.16 75.57 72.89 70.86 68.17 67.01 65.26 63.36 61.76 60.26
CEC (Zhang et al., 2021) 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28
FACT (Zhou et al., 2022a) 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94
LIMIT (Zhou et al., 2022b) 75.89 73.55 71.99 68.14 67.42 63.61 62.40 61.35 59.91 58.66 57.41
SoftNet (Kang et al., 2023) 78.07 74.58 71.37 67.54 65.37 62.60 61.07 59.37 57.53 57.21 56.75
SAVC (Song et al., 2023) 81.85 77.92 74.95 70.21 69.96 67.02 66.16 65.30 63.84 63.15 62.50
BOT (Roy et al., 2024) 82.31 78.03 75.45 70.99 71.06 67.85 67.44 66.05 64.95 64.31 63.75
SV-T (Qiu et al., 2023) SwinT 84.19 82.63 81.21 78.97 79.38 77.64 77.55 75.71 75.91 75.77 76.17
CPE-CLIP (D’Alessandro et al., 2023) CLIP-B/16 81.58 78.52 76.68 71.86 71.52 70.23 67.66 66.52 65.09 64.47 64.60
PriViLege (Park et al., 2024) CLIP-B/16 82.21 81.25 80.45 77.76 77.78 75.95 75.69 76.00 75.19 75.19 75.08
Ours ViT-B/16 88.68 86.26 85.83 83.38 83.52 81.71 81.77 81.77 81.02 80.76 81.19
Table 2: Comparison to prior works across the base and incremental sessions on miniImageNet with 60 base classes and 5-way 5-shot incremental setting.
Method PT. Backbone Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
iCaRL (Rebuffi et al., 2017) 71.77 61.85 58.12 54.60 51.49 48.47 45.90 44.19 42.71
Rebalancing (Hou et al., 2019) 72.30 66.37 61.00 56.93 53.31 49.93 46.47 44.13 42.19
TOPIC (Tao et al., 2020) 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
EEIL (Castro et al., 2018) 61.31 46.58 44.00 37.29 33.14 27.12 24.10 21.57 19.58
FSLL (Mazumder et al., 2021) 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
FSLL+SS (Mazumder et al., 2021) 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
F2M (Shi et al., 2021) 72.05 67.47 63.16 59.70 56.71 53.77 51.11 49.21 47.84
CEC (Zhang et al., 2021) 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63
MetaFSCIL (Chi et al., 2022) 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19
C-FSCIL (Hersche et al., 2022) 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41
FACT (Zhou et al., 2022a) 72.56 69.63 66.38 62.77 60.60 57.33 54.34 52.16 50.49
CLOM (Zou et al., 2022) 73.08 68.09 64.16 60.41 57.41 54.29 51.54 49.37 48.00
LIMIT (Zhou et al., 2022b) 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19
SoftNet (Kang et al., 2023) 79.77 75.08 70.59 66.93 64.00 61.00 57.81 55.81 54.68
SAVC (Song et al., 2023) 81.12 76.14 72.43 68.92 66.48 62.95 59.92 58.39 57.11
BOT (Roy et al., 2024) 84.30 79.59 75.49 71.4 68.45 65.12 62.20 60.52 59.57
SV-T (Qiu et al., 2023) SwinT 88.75 87.92 86.07 84.84 84.30 83.24 82.22 82.28 82.38
CPE-CLIP (D’Alessandro et al., 2023) CLIP-B/16 90.23 89.56 87.42 86.80 86.51 85.08 83.43 83.38 82.77
PriViLege (Park et al., 2024) CLIP-B/16 96.68 96.49 95.65 95.54 95.54 94.91 94.33 94.19 94.10
Ours ViT-B/16 97.63 97.55 97.09 97.02 97.0 96.58 96.3 96.29 96.24
4.1 Datasets and implementation details.
@@
Following existing literature on FSCIL, we evaluate CoACT on CIFAR-100 (Krizhevsky et al., 2009), CUB-200 (Wah et al., 2011), and miniImageNet (Russakovsky et al., 2015) datasets. We evaluate our new benchmark, FSCIT on a diverse set of 16 datasets, including generic object detection (Caltech101 (Fei-Fei et al., 2004), CIFAR-100 (Krizhevsky et al., 2009), CUB-200 (Wah et al., 2011), miniImageNet (Russakovsky et al., 2015), VOC 2007 (Everingham, 2008)), fine-grained recognition (OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flower102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), FGVCAircraft (Maji et al., 2013)), scene recognition (SUN397 (Xiao et al., 2010), Country211 (Radford et al., 2021)), satellite-image (EuroSAT (Helber et al., 2019), Resisc-45 (Cheng et al., 2017)), texture recognition (DTD (Cimpoi et al., 2014)), and traffic sign recognition (GTSRB (Houben et al., 2013)). By default, we divide the classes into 10 (or 9) sessions with an equal number of classes and perform 10-shot continual learning.
@@
The performance after each session is calculated as the average accuracy of all classes seen so far. Unless specified otherwise, accuracy refers to the accuracy after the last session, which is the average accuracy over all classes on the test set. We also evaluate CoACT on the traditional FSCIL setup, where the initial training is performed with a large in-distribution base session, followed by few-shot tuning over the incremental sessions. We use ViT-B/16 as the backbone for most of the experiments, while we present the detailed per-dataset results with ViT-B/32 and ViT-L/16 as well. The encoders are pre-trained on ImgeNet-21K (Russakovsky et al., 2015). We implement our framework in PyTorch and train the model using an SGD optimizer with a momentum of 0.9. The base learning rate is set to 0.1, with a batch size of 64, and the model is trained for 50 epochs for the first session and 5 epochs for the remaining sessions. A cosine LR decay scheduler is used to reduce the learning rate over the training epochs. The teacher encoder is updated with a momentum value of 0.999. For experiments with the FSCIL setup, we train the model for 25 epochs with a learning rate of 0.001. We train the model with input resolution of $224\times 224$. All other implementation details are the same as described above. All experiments are conducted with 3 random seeds, and the reported results are averaged over the three runs. All experiments are conducted on an Nvidia V100 GPU, where the training takes about 6 hours.
Table 3: Comparison to prior works across the base and incremental sessions on CIFAR-100 with 60 base classes and 5-way 5-shot incremental setting.
Method PT. Backbone Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
Rebalancing (Hou et al., 2019) 61.31 47.80 39.31 31.91 25.68 21.35 18.67 17.24 14.17
iCaRL (Rebuffi et al., 2017) 61.31 46.32 42.94 37.63 30.49 24.00 20.89 18.80 17.21
TOPIC (Tao et al., 2020) 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
IDLVQ-C (Chen & Lee, 2020) 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84
FSLL (Mazumder et al., 2021) 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
FSLL+SS (Mazumder et al., 2021) 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
F2M (Shi et al., 2021) 67.28 63.80 60.38 57.06 54.08 51.39 48.82 46.58 44.65
CEC (Zhang et al., 2021) 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14
MetaFSCIL (Chi et al., 2022) 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97
CLOM (Zou et al., 2022) 74.2 69.83 66.17 62.39 59.26 56.48 54.36 52.16 50.25
C-FSCIL (Hersche et al., 2022) 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47
LIMIT (Zhou et al., 2022b) 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23
FACT (Zhou et al., 2022a) 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10
SAVC (Song et al., 2023) 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
SoftNet (Kang et al., 2023) 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
BOT (Roy et al., 2024) 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
SV-T (Qiu et al., 2023) SwinT 86.77 82.82 80.36 77.20 76.06 74.00 72.92 71.68 69.75
CPE-CLIP (D’Alessandro et al., 2023) CLIP-B/16 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
PriViLege (Park et al., 2024) CLIP-B/16 90.88 89.39 88.97 87.55 87.83 87.35 87.53 87.15 86.06
Ours ViT-B/16 90.46 88.46 88.11 86.94 86.98 86.52 86.39 86.0 84.63
4.2 FSCIL results
In this section, we discuss the results of CoACT in the FSCIL setup. Following the existing literature, we present these results on the CIFAR-100, CUB-200, and miniImageNet datasets. In Table 1, we compare the performance of CoACT with prior works on the CUB-200 dataset, where we divide the existing methods into two groups: the first group, which trains randomly initialized models and the second group which use a pre-trained encoder. The results are reported for 100 base classes and a 10-way, 5-shot incremental learning setup. As we observe, CoACT outperforms the previous SOTA (SV-T) by 5.02%.
Results on miniImageNet are reported in Table 2, with 60 base classes and a 5-way, 5-shot incremental learning setup. As we find from this table, CoACT outperforms the prior SOTA (PriViLege) by 2.14%. Also, the performance difference increases (over the sessions) as we learn new classes, indicating that CoACT shows less forgetting than PriViLege. Finally, the results for CIFAR-100 are presented in Table 3, where PriViLege (Park et al., 2024) holds the highest result. With significantly fewer parameters (86M in ViT-B/16 compared to 149M in CLIP-B/16), our proposed CoACT achieves competitive results, performing only 1.43% below PriViLege.
Dataset Classes Sessions First Ses. class Cls/session
Caltech101 102 10 12 10
CIFAR-100 100 10 10 10
Country211 211 10 22 21
CUB-200 200 10 20 20
DTD 47 9 5 5
EuroSAT 10 10 1 1
FGVCAircraft 100 10 10 10
Food101 101 10 11 10
GTSRB 43 10 7 4
MiniImageNet 100 10 10 10
Flower102 102 10 12 10
OxfordPets 37 9 5 4
Resisc-45 45 9 5 5
StanfordCars 196 9 28 21
SUN397 397 10 45 44
VOC 2007 20 10 2 2
Table 4: Details on class splits over the continual sessions for different datasets.
@@ 4.2 FSCIL results
Caltech101
CIFAR100
Country211
CUB200
DTD
EuroSat
FGVCAircraft
Food101
GTSRB
MiniImageNet
Flowers102
OxfordPets
Resisc-45
StanfordCars
SUN397
VOC 2007
Average
CPE-CLIP 82.79 74.11 7.03 73.47 53.92 61.77 14.79 69.71 14.64 94.57 97.02 87.05 48.43 18.51 64.48 63.53 57.85
PriViLege 83.49 75.06 7.28 74.45 54.44 62.57 15.23 70.67 15.87 95.04 97.46 87.34 49.17 18.67 65.09 64.11 58.49
Pro. lear. 83.47 75.58 7.42 74.10 54.72 62.20 15.51 71.09 15.53 95.31 98.09 88.10 49.83 18.90 65.68 64.72 58.77
Lin. tun. 83.60 75.58 7.45 74.41 54.72 62.20 15.32 71.18 16.92 95.32 98.09 88.10 49.83 18.96 65.87 64.72 58.89
LoRA 83.87 75.81 7.45 74.65 54.70 62.20 15.32 71.24 17.23 95.32 98.09 88.10 49.82 18.94 65.90 64.72 58.96
CoACT 86.86 78.31 7.42 77.38 55.11 62.25 18.98 71.59 26.05 95.34 98.18 87.79 62.62 24.40 66.11 64.45 61.43
Table 5: Performance of CoACT on FSCIT across 16 datasets and its comparison to existing FSCIL methods and our baseline methods. Pro. lear. and Lin. tun. respectively refer to prototype learning and linear tuning.
4.3 FSCIT results
As the first work on FSCIT, we first establish a few baselines to better evaluate our proposed framework. To this end, we explore two of the existing SOTA from the FSCIL literature PriViLege and CPE-CLIP and also explore 3 well-known approaches from the few-shot tuning literature. First, we adopt prototype learning, which has been shown to perform well for learning incremental classes in traditional FSCIL (Song et al., 2023). To apply this approach to FSCIT, we keep the pre-trained encoder frozen during the incremental learning sessions, while the classifier ($W$) is learned with the prototypes (Song et al., 2023) of the new classes: $W=\{w_{0}^{0},w_{1}^{0},\cdot\cdot,w_{|c^{0}|}^{0}\}\cup\{w_{0}^{T},w_{1}^{T},\cdot\cdot,w_{|c^{T}|}^{T}\}$. Here, the prototype $w_{c}^{t}$ of class $c$ from session $t$ is defined as the average of the embeddings of all samples of the class: ${w}_{c}^{t}=\frac{1}{|P_{i}|}\sum_{j\in P_{i}}{f_{\theta}({{x}}_{j})}$, where $P_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\{j:y_{j}=i\}$. Next, we consider an incremental-frozen approach with linear tuning as a baseline, where we fine-tune the model only on the first session, followed by prototype learning for the remaining sessions (Zhou et al., 2022a). Finally, we use parameter-efficient fine-tuning as the third baseline, where we add LoRA to different layers of the pre-trained encoder, train them during the first setting, and perform prototype learning in incremental sessions. Additionally, to ensure extensive evaluation of the new benchmark, we conduct the experiments on 16 datasets. The classes are divided into sessions as summarized in Table 4.
In Table 5, we present the main results of our study on the FSCIT setup using the ViT-B/16 encoder. The table highlights that existing methods for FSCIL, such as CPE-CLIP and PriViLege, do not perform well on FSCIT. This is primarily because these methods heavily rely on supervised training of the base session, which is absent in the FSCIT setup, leading to reduced performance. Among the other three baselines, namely prototype tuning, linear tuning, and LoRA, performance is slightly better compared to FSCIL methods. However, CoACT surpasses all prior approaches by significant margins on all the datasets. Specifically, CoACT achieves an average improvement of 2.47% over the baselines and exhibits notable gains of up to 12.79% on individual datasets, such as Resisc-45. Additionally, we observe larger improvements on more challenging datasets. For instance, on the five datasets with the lowest accuracy (Country211, FGVCAircraft, GTSRB, Resisc-45, and StanfordCars), CoACT achieves an average improvement of 6.14% over the best-performing baseline.
@@ 4.3 FSCIT results
ViT-L/14 87.25 79.27 7.45 78.29 56.19 66.36 21.24 74.99 17.22 97.18 98.18 87.95 62.25 25.43 67.18 62.23 61.79
Swin-B 88.01 80.17 8.15 78.91 57.25 68.98 23.44 76.11 28.18 97.01 98.03 88.56 64.24 25.90 69.34 66.95 63.70
DINO ViT-B/14 87.87 79.92 7.89 79.01 57.81 68.57 22.58 76.47 30.38 96.94 98.05 88.29 63.59 25.87 69.04 65.47 63.61
CLIP ViT-B/16 (V) 84.23 78.14 6.49 72.89 55.23 68.48 24.93 72.49 27.38 95.18 97.23 86.66 57.23 23.47 64.29 62.38 61.04
Table 6: Performance of CoACT on FSCIT across 16 datasets on different encoders.
Figure 4: (left) Forgetting of learned classes in FSCIT. (right) Accuracy breakdown into the first, remaining and all sessions in FSCIT setup.
@@
Next, we report the forgetting of learned classes for each method, measured as the drop in accuracy w.r.t. the first session, in Figure 4 (left). As we observe in this figure, CoACT has the least amount of forgetting compared to the baselines, with approximately 1.5% less forgetting than both LoRA and linear tuning and 3.2% than prototype learning. We also present a breakdown of accuracies into the first incremental session, the remaining incremental sessions, and all sessions in Figure 4 (right). Here, for all methods, the first session shows particularly higher accuracy than other sessions since there is no interference (or forgetting) of other classes. While LoRA and linear tuning have higher overall accuracy than prototype tuning, the improvement mainly comes from the higher accuracy in the first session only. All baselines, however, perform relatively similarly in the remaining sessions. In contrast, CoACT shows higher improvement in both the first and remaining sessions.
Figure 5: Performance for different shots in FSCIT.
Next, we investigate the generalization of CoACT across different encoders of different sizes. Specifically, we explore ViT-B/16, ViT-B/32, ViT-L/14, Swin-B (Liu et al., 2021), DINO ViT-B/14 (Oquab et al., ), and vision encoder from CLIP ViT-B/16 (Radford et al., 2021) as the backbone. The results of this study are presented in Table 6. The results from this study show that our proposed solution generalizes across the encoder sizes. Specifically, with a larger encoder (ViT-L/14), CoACT improves the average accuracy to 61.79%. On individual datasets, we find up to 3.4% improvements over the default ViT-B/16 encoder. Similarly, CoACT shows strong performances on the Swin-B, DINO ViT-B/14, and CLIP ViT-B/16 encoders.